# GLGENN: A Novel Parameter-Light Equivariant Neural Networks Architecture Based on Clifford Geometric Algebras

### v1

@@ Appendix A Notation
$C\kern-0.90005pt\ell_{p,q,r}$ (Clifford) Geometric algebra (GA) over the real $\mathbb{R}^{p,q,r}$ or complex $\mathbb{C}^{p+q,0,r}$ vector space page 2.1
$C\kern-0.90005pt\ell_{p,q}$ Non-degenerate geometric algebra $C\kern-0.90005pt\ell_{p,q,0}$ page 2.1
$\Lambda_{r}$ Grassmann subalgebra $C\kern-0.90005pt\ell_{0,0,r}$ of $C\kern-0.90005pt\ell_{p,q,r}$ page 2.1
@@
$e_{1},\ldots,e_{n}$ Generators of $C\kern-0.90005pt\ell_{p,q,r}$ page 1
$C\kern-0.90005pt\ell^{0}$ Subspace of grade 00 page 1
$C\kern-0.90005pt\ell^{k}_{p,q,r}$ Subspace of grade $k$, where $k=1,\ldots,n$ page 1
$\{C\kern-0.90005pt\ell^{k}_{p,q}{\rm\Lambda}^{l}_{r}\}$ Subspace of $C\kern-0.90005pt\ell_{p,q,r}$ spanned by the elements of the form $ab$, where $a\in C\kern-0.90005pt\ell^{k}_{p,q}$ and $b\in{\rm\Lambda}^{l}_{r}$ page 75
@@
$\widehat{U}$ Grade involute of $U\in C\kern-0.90005pt\ell_{p,q,r}$ formula (2)
$\widetilde{U}$ Reversion of $U\in C\kern-0.90005pt\ell_{p,q,r}$ formula (2)
$\widehat{\widetilde{U}}$ Clifford conjugate of $U\in C\kern-0.90005pt\ell_{p,q,r}$ page 2
$C\kern-0.90005pt\ell^{(0)}_{p,q,r}$, $C\kern-0.90005pt\ell^{(1)}_{p,q,r}$ Even and odd subspaces of $C\kern-0.90005pt\ell_{p,q,r}$ formula (3)
$(C\kern-0.90005pt\ell^{(0)\times}_{p,q,r}\cup C\kern-0.90005pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ Product of two groups: $\{ab\;|\;a\in C\kern-0.90005pt\ell^{(0)\times}_{p,q,r}\cup C\kern-0.90005pt\ell^{(1)\times}_{p,q,r},\;b\in\Lambda^{\times}_{r}\}$ Lemma 2.1
$\langle\rangle_{(l)}$ Projection onto the subspace $C\kern-0.90005pt\ell^{(l)}_{p,q,r}$, where $l=0,1$ page 3
$C\kern-0.90005pt\ell^{\overline{m}}_{p,q,r}$ Subspaces determined by grade involution and reversion formula (4)
$\langle\rangle_{\overline{m}}$ Projection onto the subspace $C\kern-0.90005pt\ell^{\overline{m}}_{p,q,r}$, where $m=0,1,2,3$ page 4
@@
$\tilde{{\rm ad}}^{1}$ $\tilde{{\rm ad}}$ restricted to $C\kern-0.90005pt\ell^{1}_{p,q,r}$ page 34
@@
$\psi$, $\chi$ Norm functions of $C\kern-0.90005pt\ell_{p,q,r}$ elements formula (11)
@@
$\Gamma^{\overline{k}}_{p,q,r}={\rm Q}^{\overline{k}}_{p,q,r}$, Groups preserving the subspaces $C\kern-0.90005pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, under ${\rm ad}$ formulas (16) and (18)–(19)
$\check{\Gamma}^{\overline{k}}_{p,q,r}=\check{{\rm Q}}^{\overline{k}}_{p,q,r}$ Groups preserving the subspaces $C\kern-0.90005pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, under $\check{{\rm ad}}$ formulas (17) and (20)–(21)
${\rm Z}_{p,q,r}$ Center of $C\kern-0.90005pt\ell_{p,q,r}$ formula (72)
${\rm Z}^{k}_{p,q,r}$, $\check{{\rm Z}}^{k}_{p,q,r}$ Centralizers and twisted centralizers respectively of $C\kern-0.90005pt\ell^{k}_{p,q,r}$, $k=0,\ldots,n$ page 21
${\rm Z}^{\overline{k}}_{p,q,r}$, $\check{{\rm Z}}^{\overline{k}}_{p,q,r}$ Centralizers and twisted centralizers respectively of $C\kern-0.90005pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$ formulas (103)–(104)
@@
In this section, we provide an example of the key theoretical concepts and constructions considered in Section 2. Let us consider the Minkowski space $V=\mathbb{R}^{1,3}$ equipped with the Minkowski metric $\eta=\mathrm{diag}(1,-1,-1,-1)$. In this case, we have $p=1$, $q=3$, $r=0$, and $n=p+q=4$. The corresponding geometric (Clifford) algebra is $C\kern-1.00006pt\ell_{1,3}=C\kern-1.00006pt\ell(\mathbb{R}^{1,3})$, which is the real non-degenerate $162^{4}=$ 16-dimensional algebra with the identity element $e$ and generators $e_{1},e_{2},e_{3},e_{4}$ that satisfy:
@@
which implies that the distinct generators anticommute, the first generator squared gives $+e$, and the other three generators squared give $-e$:
@@
which we call the subspaces of scalars, vectors, bivectors, trivectors, and $4$-vectors respectively. An arbitrary element $U\in C\kern-1.00006pt\ell_{1,3}$ can be written as
@@
and for any other associative algebra $A$ over $V$ with the unity $e_{A}$ and any linear map $k:V\rightarrow A$, such that $k^{2}(x)=\mathfrak{q}(x)e_{A}$ for any $x\in V$, there exists a unique algebra homomorphism $f:C\kern-1.00006pt\ell(V)\rightarrow A$, such that $f\circ j=k$, i.e. the corresponding diagram is commutative:
${V}$${C\kern-1.00006pt\ell(V)}$${A}$$\scriptstyle{j}$$\scriptstyle{k}$$\scriptstyle{f}$
@@
Let us consider the opposite algebra $C\kern-1.00006pt\ell(V)^{o}$, in which the product of $x$ and $y$ is given by $yx$. Then there exists a unique algebra homomorphism $\tau:C\kern-1.00006pt\ell(V)\rightarrow C\kern-1.00006pt\ell(V)^{o}$ as in the diagram below
${V}$${C\kern-1.00006pt\ell(V)}$${C\kern-1.00006pt\ell(V)^{o}}$$\scriptstyle{j}$$\scriptstyle{j}$$\scriptstyle{\tau}$
@@
${V}$${C\kern-1.00006pt\ell(V)}$${C\kern-1.00006pt\ell(V)}$$\scriptstyle{j}$$\scriptstyle{\alpha_{0}}$$\scriptstyle{\alpha}$
@@ Appendix D Centralizers and Twisted Centralizers
The paper Filimoshina & Shirokov, 2024b studies explicit forms of the sets ${{\rm Z}}^{m}_{p,q,r}$ and $\check{{\rm Z}}^{m}_{p,q,r}$ in the case of arbitrary $m=0,\ldots,n$. In Remark D.1, for the readers’ convenience, we write out several particular cases of these sets, which we further use. A few words about the notation in this remark and below. The spaces $C\kern-1.00006pt\ell^{k}_{p,q}$ and ${\rm\Lambda}^{k}_{r}$, $k=0,\ldots,n$, are regarded as subspaces of $C\kern-1.00006pt\ell_{p,q,r}$. By $\{C\kern-1.00006pt\ell^{k}_{p,q}{\rm\Lambda}^{l}_{r}\}$, we denote the subspace of $C\kern-1.00006pt\ell_{p,q,r}$ spanned by the elements of the form $ab$, where $a\in C\kern-1.00006pt\ell^{k}_{p,q}$ and $b\in{\rm\Lambda}^{l}_{r}$.
@@
$\langle{\rm Z}^{4}_{p,q,r}\rangle_{(0)}\!=\!\left\{\begin{array}[]{lll}\Lambda^{(0)}_{r}\oplus\{C\kern-1.00006pt\ell^{1}_{p,q}\Lambda^{n-2}_{r}\}\oplus\{C\kern-1.00006pt\ell^{2}_{p,q}\Lambda^{n-3}_{r}\},&&\mbox{$n$ is oddor $r=n$;}\\ \Lambda^{(0)}_{r}\oplus\{C\kern-1.00006pt\ell^{1}_{p,q}\Lambda^{n-3}_{r}\}\oplus\{C\kern-1.00006pt\ell^{2}_{p,q}\Lambda^{n-4}_{r}\}\oplus C\kern-1.00006pt\ell^{n}_{p,q,r},&&\mbox{$n$ is even, $r\neq n$}.\end{array}\right.$ (108)
@@ Appendix E Ordinary Lipschitz Groups, Spin Groups, Adjoint and Twisted Adjoint Representations
In the non-degenerate geometric algebras $C\kern-1.00006pt\ell_{p,q}$, the Lipschitz group is defined as the group of all invertible elements preserving the grade-$1$ subspace under the twisted adjoint representation $\check{{\rm ad}}$. Generalizing this definition to the case of the degenerate $C\kern-1.00006pt\ell_{p,q,r}$, we similarly define the Lipschitz group $\tilde{\Gamma}^{1}_{p,q,r}$ as
@@
Also let us consider the twisted adjoint representation, which is introduced by Atiyah, Bott, and Shapiro in Atiyah et al., 1964. In this paper, the twisted adjoint representation is defined in the case of the non-degenerate algebra $C\kern-1.00006pt\ell_{0,q,0}$. It acts on the Lipschitz group $\tilde{\Gamma}^{1}_{0,q,0}$ (113) in the way $\check{{\rm ad}}:\Gamma^{\pm}_{0,q,0}\rightarrow{\rm Aut}(C\kern-1.00006pt\ell^{1}_{0,q,0})$ as $T\rightarrow\check{{\rm ad}}_{T}$, where $\check{{\rm ad}}_{T}:C\kern-1.00006pt\ell^{1}_{0,q,0}\rightarrow C\kern-1.00006pt\ell^{1}_{0,q,0}$ is defined for elements of the grade-$1$ subspace (vectors) as
@@
Let us prove $\ker(\tilde{{\rm ad}})\subseteq\Lambda^{\times}_{r}$. Suppose $T\in\ker(\tilde{{\rm ad}})$; then $\widehat{T}U_{1}T^{-1}=U_{1}$ for any $U_{1}\in C\kern-1.00006pt\ell^{1}_{p,q,r}$. Thus, $T\in\Lambda^{\times}_{r}$ by the statement (88) of Remark D.1.
Now we must only prove that $\Lambda^{\times}_{r}\subseteq\ker(\tilde{{\rm ad}})$. Suppose $T\in\Lambda^{\times}_{r}$; then $TC\kern-1.00006pt\ell^{(0)}_{p,q,r}=C\kern-1.00006pt\ell^{(0)}_{p,q,r}T$ and $\widehat{T}C\kern-1.00006pt\ell^{1}_{p,q,r}=C\kern-1.00006pt\ell^{1}_{p,q,r}T$ by Lemma D.3 and the statement (88) of Remark D.1 respectively. Since any odd basis element can be represented as a product of an odd number of generators, we obtain $\widehat{T}C\kern-1.00006pt\ell^{(1)}_{p,q,r}=C\kern-1.00006pt\ell^{(1)}_{p,q,r}T$. Thus, $TU_{(0)}T^{-1}+\widehat{T}U_{(1)}T^{-1}=U$ for all $U=U_{(0)}+U_{(1)}\in C\kern-1.00006pt\ell_{p,q,r}$, where $U_{(0)}\in C\kern-1.00006pt\ell^{(0)}_{p,q,r}$ and $U_{(1)}\in C\kern-1.00006pt\ell^{(1)}_{p,q,r}$, and the proof is completed. ∎
@@
and get $\widehat{T^{-1}}T\in\Lambda^{\times}_{r}$, which implies $T^{-1}\widehat{T}=\widehat{\widehat{T^{-1}}T}\in\Lambda^{\times}_{r}$. Since $\Lambda^{\times}_{r}=\ker(\tilde{{\rm ad}})$ by Lemma E.1, we obtain that $(T^{-1}\widehat{T})X=X(T^{-1}\widehat{T})$ for any even $X\in C\kern-1.00006pt\ell^{(0)}_{p,q,r}$ and $(T^{-1}\widehat{T})X=X\widehat{(T^{-1}\widehat{T})}$ for any odd $X\in C\kern-1.00006pt\ell^{(1)}_{p,q,r}$. Therefore,
@@
[…]
@@
$\check{{\rm Q}}^{\overline{1}}_{p,q,r}:=\{T\in C\kern-1.00006pt\ell^{\times}_{p,q,r}:\quad\widetilde{T}T\in\check{{\rm Z}}^{1\times}_{p,q,r}=\ker(\tilde{{\rm ad}}),\quad\widehat{\widetilde{T}}T\in\check{{\rm Z}}^{1\times}_{p,q,r}=\ker(\tilde{{\rm ad}})\},$ (157)
@@
where $\ker(\tilde{{\rm ad}})$ (130) is the kernel of the twisted adjoint representation $\tilde{{\rm ad}}$ (123), the sets $\check{{\rm Z}}^{2}_{p,q,r}$ (92) and $\check{{\rm Z}}^{3}_{p,q,r}$ (93) are the twisted centralizers of the subspaces $C\kern-1.00006pt\ell^{2}_{p,q,r}$ and $C\kern-1.00006pt\ell^{3}_{p,q,r}$ respectively, and ${\rm Z}^{4}_{p,q,r}$ (Remark D.1) is the centralizer of the subspace $C\kern-1.00006pt\ell^{4}_{p,q,r}$ (see Appendix D). We use $\check{\;}$ in the notation of the groups (157)–(160) due to Theorem F.2 below. We have (see (130)):
@@
[…]
@@
[…]
@@
Suppose $T\in\tilde{\Gamma}^{\overline{1}}_{p,q,r}$; then $\widetilde{T}T=W\in\ker(\tilde{{\rm ad}})$ and $\widehat{\widetilde{T}}T=U\in\ker(\tilde{{\rm ad}})$. Therefore, $T=\widetilde{T^{-1}}W$ and $\widehat{T^{-1}}=\widehat{U^{-1}}\widetilde{T}$. Thus, $\widehat{T^{-1}}T=(\widehat{U^{-1}}\widetilde{T})(\widetilde{T^{-1}}W)=\widehat{U^{-1}}W\in\ker(\tilde{{\rm ad}})=\Lambda^{\times}_{r}$. Now we apply Theorem 4.7 (Filimoshina & Shirokov, 2023):

### v2

@@ Appendix A Notation
$C\kern-0.92505pt\ell_{p,q,r}$ (Clifford) Geometric algebra (GA) over the real $\mathbb{R}^{p,q,r}$ or complex $\mathbb{C}^{p+q,0,r}$ vector space page 2.1
$C\kern-0.92505pt\ell_{p,q}$ Non-degenerate geometric algebra $C\kern-0.92505pt\ell_{p,q,0}$ page 2.1
$\Lambda_{r}$ Grassmann subalgebra $C\kern-0.92505pt\ell_{0,0,r}$ of $C\kern-0.92505pt\ell_{p,q,r}$ page 2.1
@@
$e_{1},\ldots,e_{n}$ Generators of $C\kern-0.92505pt\ell_{p,q,r}$ page 1
$C\kern-0.92505pt\ell^{0}$ Subspace of grade 0 page 1
$C\kern-0.92505pt\ell^{k}_{p,q,r}$ Subspace of grade $k$, where $k=1,\ldots,n$ page 1
$\{C\kern-0.92505pt\ell^{k}_{p,q}{\rm\Lambda}^{l}_{r}\}$ Subspace of $C\kern-0.92505pt\ell_{p,q,r}$ spanned by the elements of the form abab, where $a\in C\kern-0.92505pt\ell^{k}_{p,q}$ and $b\in{\rm\Lambda}^{l}_{r}$ page 75
@@
$\widehat{U}$ Grade involute of $U\in C\kern-0.92505pt\ell_{p,q,r}$ formula (2)
$\widetilde{U}$ Reversion of $U\in C\kern-0.92505pt\ell_{p,q,r}$ formula (2)
$\widehat{\widetilde{U}}$ Clifford conjugate of $U\in C\kern-0.92505pt\ell_{p,q,r}$ page 2
$C\kern-0.92505pt\ell^{(0)}_{p,q,r}$, $C\kern-0.92505pt\ell^{(1)}_{p,q,r}$ Even and odd subspaces of $C\kern-0.92505pt\ell_{p,q,r}$ formula (3)
$(C\kern-0.92505pt\ell^{(0)\times}_{p,q,r}\cup C\kern-0.92505pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ Product of two groups: $\{ab\;|\;a\in C\kern-0.92505pt\ell^{(0)\times}_{p,q,r}\cup C\kern-0.92505pt\ell^{(1)\times}_{p,q,r},\;b\in\Lambda^{\times}_{r}\}$ Lemma 2.1
$\langle\rangle_{(l)}$ Projection onto the subspace $C\kern-0.92505pt\ell^{(l)}_{p,q,r}$, where $l=0,1$ page 3
$C\kern-0.92505pt\ell^{\overline{m}}_{p,q,r}$ Subspaces determined by grade involution and reversion formula (4)
$\langle\rangle_{\overline{m}}$ Projection onto the subspace $C\kern-0.92505pt\ell^{\overline{m}}_{p,q,r}$, where $m=0,1,2,3$ page 4
@@
$\tilde{{\rm ad}}^{1}$ $\tilde{{\rm ad}}$ restricted to $C\kern-0.92505pt\ell^{1}_{p,q,r}$ page 34
@@
$\psi$, $\chi$ Norm functions of $C\kern-0.92505pt\ell_{p,q,r}$ elements formula (11)
@@
$\Gamma^{\overline{k}}_{p,q,r}={\rm Q}^{\overline{k}}_{p,q,r}$, Groups preserving the subspaces $C\kern-0.92505pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, under $\rm ad$} formulas (16) and (18)–(19)
$\check{\Gamma}^{\overline{k}}_{p,q,r}=\check{{\rm Q}}^{\overline{k}}_{p,q,r}$ Groups preserving the subspaces $C\kern-0.92505pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, under $\check{{\rm ad}}$ formulas (17) and (20)–(21)
${\rm Z}_{p,q,r}$ Center of $C\kern-0.92505pt\ell_{p,q,r}$ formula (72)
${\rm Z}^{k}_{p,q,r}$, $\check{{\rm Z}}^{k}_{p,q,r}$ Centralizers and twisted centralizers respectively of $C\kern-0.92505pt\ell^{k}_{p,q,r}$, $k=0,\ldots,n$ page 21
${\rm Z}^{\overline{k}}_{p,q,r}$, $\check{{\rm Z}}^{\overline{k}}_{p,q,r}$ Centralizers and twisted centralizers respectively of $C\kern-0.92505pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$ formulas (103)–(104)
@@
In this section, we provide an example of the key theoretical concepts and constructions considered in Section 2. Let us consider the Minkowski space $V=\mathbb{R}^{1,3}$ equipped with the Minkowski metric $\eta=\mathrm{diag}(1,-1,-1,-1)$. In this case, we have $p=1$, $q=3$, $r=0$, and $n=p+q=4$. The corresponding geometric (Clifford) algebra is $C\kern-1.00006pt\ell_{1,3}=C\kern-1.00006pt\ell(\mathbb{R}^{1,3})$, which is the real non-degenerate $2^{4}=16$-dimensional algebra with the identity element $e$ and generators $e_{1},e_{2},e_{3},e_{4}$ that satisfy:
@@
which implies that the distinct generators anticommute, the first generator squared gives +e+e, and the other three generators squared give e-e:
@@
which we call the subspaces of scalars, vectors, bivectors, trivectors, and 44-vectors respectively. An arbitrary element $U\in C\kern-1.00006pt\ell_{1,3}$ can be written as
@@
and for any other associative algebra AA over $V$ with the unity $e_{A}$ and any linear map $k:V\rightarrow A$, such that $k^{2}(x)=\mathfrak{q}(x)e_{A}$ for any $x\in V$, there exists a unique algebra homomorphism $f:C\kern-1.00006pt\ell(V)\rightarrow A$, such that $f\circ j=k$, i.e. the corresponding diagram is commutative:
[…]
@@
Let us consider the opposite algebra $C\kern-1.00006pt\ell(V)^{o}$, in which the product of $x$ and $y$ is given by yxyx. Then there exists a unique algebra homomorphism $\tau:C\kern-1.00006pt\ell(V)\rightarrow C\kern-1.00006pt\ell(V)^{o}$ as in the diagram below
[…]
@@
[…]
@@ Appendix D Centralizers and Twisted Centralizers
The paper Filimoshina & Shirokov, 2024b studies explicit forms of the sets ${{\rm Z}}^{m}_{p,q,r}$ and $\check{{\rm Z}}^{m}_{p,q,r}$ in the case of arbitrary $m=0,\ldots,n$. In Remark D.1, for the readers’ convenience, we write out several particular cases of these sets, which we further use. A few words about the notation in this remark and below. The spaces $C\kern-1.00006pt\ell^{k}_{p,q}$ and ${\rm\Lambda}^{k}_{r}$, $k=0,\ldots,n$, are regarded as subspaces of $C\kern-1.00006pt\ell_{p,q,r}$. By $\{C\kern-1.00006pt\ell^{k}_{p,q}{\rm\Lambda}^{l}_{r}\}$, we denote the subspace of $C\kern-1.00006pt\ell_{p,q,r}$ spanned by the elements of the form abab, where $a\in C\kern-1.00006pt\ell^{k}_{p,q}$ and $b\in{\rm\Lambda}^{l}_{r}$.
@@
$\langle{\rm Z}^{4}_{p,q,r}\rangle_{(0)}\!=\!\left\{\begin{array}[]{lll}\Lambda^{(0)}_{r}\oplus\{C\kern-1.00006pt\ell^{1}_{p,q}\Lambda^{n-2}_{r}\}\oplus\{C\kern-1.00006pt\ell^{2}_{p,q}\Lambda^{n-3}_{r}\},\quad&&\mbox{$n$ is odd or $r=n$;}\\ \Lambda^{(0)}_{r}\oplus\{C\kern-1.00006pt\ell^{1}_{p,q}\Lambda^{n-3}_{r}\}\oplus\{C\kern-1.00006pt\ell^{2}_{p,q}\Lambda^{n-4}_{r}\}\oplus C\kern-1.00006pt\ell^{n}_{p,q,r},\quad&&\mbox{$n$ is even, $r\neq n$}.\end{array}\right$. (108)
@@ Appendix E Ordinary Lipschitz Groups, Spin Groups, Adjoint and Twisted Adjoint Representations
In the non-degenerate geometric algebras $C\kern-1.00006pt\ell_{p,q}$, the Lipschitz group is defined as the group of all invertible elements preserving the grade-11 subspace under the twisted adjoint representation $\check{{\rm ad}}$. Generalizing this definition to the case of the degenerate $C\kern-1.00006pt\ell_{p,q,r}$, we similarly define the Lipschitz group $\tilde{\Gamma}^{1}_{p,q,r}$ as
@@
Also let us consider the twisted adjoint representation, which is introduced by Atiyah, Bott, and Shapiro in Atiyah et al., 1964. In this paper, the twisted adjoint representation is defined in the case of the non-degenerate algebra $C\kern-1.00006pt\ell_{0,q,0}$. It acts on the Lipschitz group $\tilde{\Gamma}^{1}_{0,q,0}$ (113) in the way $\check{{\rm ad}}:\Gamma^{\pm}_{0,q,0}\rightarrow{\rm Aut}(C\kern-1.00006pt\ell^{1}_{0,q,0})$ as $T\rightarrow\check{{\rm ad}}_{T}$, where $\check{{\rm ad}}_{T}:C\kern-1.00006pt\ell^{1}_{0,q,0}\rightarrow C\kern-1.00006pt\ell^{1}_{0,q,0}$ is defined for elements of the grade-11 subspace (vectors) as
@@
[…]
@@
[…]
@@
${\rm Q}^{\overline{1}}_{p,q}={\rm Q}^{\overline{3}}_{p,q}={\rm Q},\qquad{\rm Q}^{\overline{0}}_{p,q}={\rm Q}^{\overline{2}}_{p,q}=\left\{\begin{array}[]{lll}{\rm Q},&&n=1,2,3{\rm\;mod\;}{4},\\ {\rm Q}^{\prime},&&n=0{\rm\;mod\;}{4},\end{array}\right.\quad n\neq 2,3,$
@@
[…]
@@
[…]
@@
$\check{{\rm Q}}^{\overline{1}}_{p,q}=\check{{\rm Q}}^{\overline{3}}_{p,q}={\rm Q}^{\pm},\qquad\check{{\rm Q}}^{\overline{0}}_{p,q}=\check{{\rm Q}}^{\overline{2}}_{p,q}=\left\{\begin{array}[]{lll}{\rm Q}^{\pm},&&n=1,2,3{\rm\;mod\;}{4},\\ {\rm Q}^{\prime},&&n=0{\rm\;mod\;}{4},\end{array}\right.\;n\neq 1,2,$
@@
We get $\check{{\rm Q}}^{\overline{1}}_{p,q,r}\subseteq\check{{\rm Q}}^{\overline{3}}_{p,q,r}$, using $\check{{\rm Z}}^{1}_{p,q,r}\subseteq\check{{\rm Z}}_{p,q,r}^{3}$ (see Remark D.1). For fixed $k=0,1,2,3{\rm\;mod\;}{4}$ and $m=(k-1){\rm\;mod\;}{4}$, $l=(k+1){\rm\;mod\;}{4}$, we obtain
@@
[…]